### v2

# Geometric Rectification of Surface Activity Induced Flow in Confined Channels

Abstract
Conventional pressure-driven flow obeys Poiseuille’s law, with the mean velocity scaling as $u\propto r^{2}$ under confinement. Here we identify a distinct transport mode driven by surface activity (e.g., mass exchange or boundary slip gradients), which is rectified by geometric asymmetry into a net axial flux. Using a minimal exchange model, we show that this mechanism exhibits four defining signatures that contrast sharply with pressure-driven Poiseuille flow: (i) an inverted confinement scaling $u\propto r^{-1}$; (ii) leading-order viscosity independence for prescribed $s$; (iii) macroscopic length amplification ($Q\propto L$); and (iv) linear superposition with pressure-driven flows. These results establish confined channels as active geometric rectifiers and provide a unified framework for surface-induced transport from microfluidic to biological settings.
Transport in confined channels is classically constrained by Poiseuille scaling: for fixed pressure drop $\Delta P$ and length $L$, the mean velocity scales as $u\propto r^{2}$ and therefore decreases rapidly as the channel narrows. Recent experiments identified a transport mode that defies this paradigm: distributed volume exchange across the wall can be rectified into net axial flow, even with zero applied pressure drop [Li2020]. Crucially, this exchange-driven transport strengthens under confinement, following the inverted scaling $u\propto r^{-1}$ (“narrower-is-faster”). Such an inverted confinement scaling is a strong diagnostic of surface-induced driving: it points to a mechanism whose strength scales with perimeter rather than cross-sectional area [Li2020]. More generally, spatially heterogeneous surface activity creates spatial gradients in thermodynamic potentials; such gradients constitute forces that drive fluid motion, which confinement and geometry can rectify into a net axial transport. In the present work, we focus on wall-normal mass exchange as a minimal and experimentally relevant realization of such surface activity.
While suction/injection flows in porous-walled channels are classical and are typically studied under an imposed pressure drop [Berman1953, Sellars1955, Yuan1956, Oxarango2004, Galowin1974], a minimal framework for the pressure-free, zero-net-exchange transport branch remains unclear.
How can a system, even in the absence of both an imposed pressure drop and any net mass exchange, organize distributed local exchange into a unidirectional current?
In this Letter, we derive a minimal hydrodynamic theory for this phenomenon and show that directed transport is symmetry controlled: a net current arises only when the combined exchange-and-geometry configuration lacks mirror symmetry, analogous to ratchets and Brownian motors where unbiased driving is rectified by broken symmetry [Reimann2002, Marchesoni2009].
We consider an incompressible Newtonian fluid in a slender channel of length $L$ and varying cross-sectional area $A(x)$. The walls are exchange active, with local volumetric source rate per unit length $s(x)$ (positive for injection, negative for removal). We consider the slender-channel limit, $\varepsilon\equiv a/L\ll 1$, where $a$ is the characteristic transverse scale (e.g., the radius $r$ for a cylindrical tube). In this regime, the flow is governed by leading-order lubrication theory. Define the volumetric flux $q(x)=A(x)u(x)$, where $u(x)$ is the cross-section-averaged axial velocity. Mass conservation gives
$\frac{\mathrm{d}q}{\mathrm{d}x}=s(x),$ (1)
and we focus on the zero-net-exchange regime,
$\int_{0}^{L}s(x)\,\mathrm{d}x=0,$ (2)
which ensures the channel neither accumulates nor depletes fluid globally (schematic in Fig. 1).
Figure 1: Exchange-driven transport model. (a) Schematic of a confined channel with varying width. The walls exhibit distributed injection ($s>0$) and removal ($s<0$). (b) Geometry-dependent sensitivity profile $\phi(x)$, which weights the contribution of local wall exchange to the net axial flux. In a uniform channel (dashed), $\phi(x)$ decreases linearly along the channel, whereas in a widening channel (solid) it is biased toward the narrow end, so that exchange occurring upstream contributes more strongly to the net flow.
Integrating Eq. (1) yields
$q(x)=Q+\int_{0}^{x}s(x^{\prime})\,\mathrm{d}x^{\prime},$ (3)
where $Q$ is the unknown constant throughput.
To determine $Q$, we close the model with the leading-order lubrication response,
$q(x)=-\kappa(x)\,\frac{\mathrm{d}P}{\mathrm{d}x},$ (4)
where $\kappa(x)>0$ is the local hydraulic conductance, which depends on the cross-sectional geometry and fluid viscosity (for a circular tube of radius $r(x)$, $\kappa(x)=\pi r(x)^{4}/(8\mu)$.)
We validated this one-dimensional reduction against full two-dimensional Stokes finite-element simulations, finding that the relative error scales as $O(\varepsilon^{2})$ with slenderness $\varepsilon$ (Fig. 2).
Figure 2: Validation of lubrication theory against 2D finite-element simulations. Relative error between 2D FEM and the 1D lubrication prediction versus slenderness $\varepsilon$, consistent with the expected $O(\varepsilon^{2})$ accuracy.
Define the imposed pressure drop by
$\Delta P\equiv P(0)-P(L).$ (5)
Using $\mathrm{d}P/\mathrm{d}x=-q/\kappa$ gives
$\Delta P=\int_{0}^{L}\frac{q(x)}{\kappa(x)}\,\mathrm{d}x.$ (6)
Define the total hydraulic resistance,
$\mathcal{R}=\int_{0}^{L}\frac{\mathrm{d}x}{\kappa(x)}.$ (7)
Substituting Eq. (3) into Eq. (6) and solving for the throughput yields
$Q=Q_{\mathrm{press}}+Q_{\mathrm{ex}},$ (8)
with
$Q_{\mathrm{press}}\equiv\frac{\Delta P}{\mathcal{R}},$
$Q_{\mathrm{ex}}\equiv-\frac{1}{\mathcal{R}}\int_{0}^{L}\frac{1}{\kappa(x)}\left[\int_{0}^{x}s(x^{\prime})\,\mathrm{d}x^{\prime}\right]\mathrm{d}x.$
This decomposition makes explicit that the net throughput is the linear superposition of a pressure-driven contribution and an exchange-driven contribution.
In this Letter, we focus on the pressure-free branch $\Delta P=0$, for which $Q=Q_{\mathrm{ex}}$ and transport arises solely from distributed wall exchange. To reveal the rectification mechanism in this branch, we rewrite the exchange-driven contribution $Q_{\mathrm{ex}}$ by exchanging the order of integration, obtaining the projection form
$Q=-\int_{0}^{L}s(x)\,\phi(x)\,\mathrm{d}x,$ (9)
with the geometry-dependent sensitivity profile
$\phi(x)=\frac{1}{\mathcal{R}}\int_{x}^{L}\frac{\mathrm{d}\xi}{\kappa(\xi)}.$ (10)
Here $\phi(x)$ is a geometry-dependent sensitivity that quantifies how strongly wall exchange at position $x$ contributes to the net through-flow. It is dimensionless, bounded ($0\leq\phi\leq 1$), and equals the fraction of the total hydraulic resistance lying downstream of $x$. As a result, geometric asymmetry renders $\phi(x)$ spatially non-uniform, so injection and suction with zero net exchange contribute unequally, yielding rectification already in the linear Stokes regime.
The exchange-driven throughput is viscosity independent at leading order. This follows directly from the rectification law, Eq. (9). That law expresses the net flow as the projection of the wall exchange profile $s(x)$ onto a geometry-dependent sensitivity profile $\phi(x)$. Because the viscosity $\mu$ enters both the local conductance $\kappa(x)$ and the total hydraulic resistance $\mathcal{R}$, it cancels in the ratio defining $\phi(x)$. Mathematically, this formulation is adjoint to the pressure-driven Stokes problem and closely related to reciprocal-theorem approaches for boundary-driven pumping [MichelinLauga2015, MasoudStone2019, HappelBrenner1983].
Furthermore, Equation (9) reveals that net transport is symmetry controlled. If the combined configuration is mirror symmetric about $L/2$, with $\kappa(x)=\kappa(L-x)$ and s(x)=s(Lx)s(x)=s(L-x), then $\phi(x)+\phi(L-x)=1$ and, using the zero-net-exchange constraint Eq. (2), the contributions cancel, giving $Q=0$. Directed transport therefore requires broken mirror symmetry in the exchange profile, the geometry, or their combination (Fig. 3). Equivalently, unbiased wall exchange produces net transport only if the combined configuration lacks mirror symmetry.
Figure 3: Mechanisms of hydrodynamic rectification. (I) Symmetric null: in a uniform channel, mirror-symmetric exchange profiles cancel, yielding $Q=0$. (II) Exchange rectification: breaking symmetry in $s(x)$ produces a net flux. (III) Geometric rectification: in a nonuniform channel, a symmetric $s(x)$ is weighted by an asymmetric sensitivity profile $\phi(x)$, giving $Q\neq 0$.
We derive a scaling bound consistent with experiments [Li2020]. Using Eq. (2), we rewrite Eq. (9) as
$Q=-\int_{0}^{L}s(x)\,[\phi(x)-\bar{\phi}]\,\mathrm{d}x,\qquad\bar{\phi}\equiv L^{-1}\int_{0}^{L}\phi(x)\,\mathrm{d}x.$
If $|s(x)|\leq S_{0}$, then
$|Q|$ $\leq\int_{0}^{L}|s(x)|\,|\phi(x)-\bar{\phi}|\,\mathrm{d}x$ (11)
$\leq S_{0}\int_{0}^{L}|\phi(x)-\bar{\phi}|\,\mathrm{d}x$
$\leq\frac{1}{2}\,S_{0}L,$
since $0\leq\phi\leq 1$ implies $\int_{0}^{L}|\phi-\bar{\phi}|\,\mathrm{d}x\leq L/2$. Thus $|Q|=O(S_{0}L)$. This scaling is notable: exchange-driven throughput grows with $L$, whereas pressure-driven throughput decays as $1/L$ at fixed $\Delta P$.
For a uniform channel where $\phi(x)=1-x/L$, the optimal zero-mean bounded activity consists of upstream injection and downstream removal (or vice versa),
$s_{\mathrm{opt}}(x)=S_{0}\,\operatorname{sgn}(L/2-x),$ (12)
yielding the explicit maximum $|Q|=S_{0}L/4$. This provides a simple design rule for exchange-driven pumping: at fixed bounds, separate source and sink regions as far as possible to minimize internal cancellation.
Comparing exchange-driven transport ($|Q_{\mathrm{ex}}|\sim S_{0}L$) with pressure-driven transport ($|Q_{\mathrm{press}}|\sim\kappa\Delta P/L$) defines a regime boundary. For a cylindrical tube, $\kappa\sim r^{4}/\mu$ and a typical exchange scale $S_{0}\sim 2\pi rv_{w}$ (with wall-normal exchange speed $v_{w}$), giving the ratio $\Gamma\sim\mu v_{w}L^{2}/(r^{3}\Delta P)$ (up to constants) and the crossover radius,
$r_{c}\sim\left(\frac{\mu v_{w}L^{2}}{\Delta P}\right)^{1/3}.$ (13)
For $r<r_{c}$, the system enters a surface-dominated regime where exchange-driven rectification can compete with (or exceed) pressure-driven transport (Fig. 4).
Figure 4: Optimization and scaling. (a) In a uniform channel, throughput is maximized by concentrating injection and removal toward opposite ends. (b) Comparison of scaling laws. Pressure-driven flow (red) weakens in tight confinement ($u\propto r^{2}$), while exchange-driven flow (blue) accelerates ($u\propto r^{-1}$), dominating below $r_{c}$.
While derived in 1D, this formulation is the exact leading-order description of Stokes flows in slender 2D and 3D conduits. Integrating $\nabla\cdot\mathbf{u}=0$ over a cross-section recovers Eq. (1), where $s(x)$ is the net wall-normal flux per unit axial length. In 3D this is a line integral around the cross-section boundary,
$s(x)=\oint_{\partial A(x)}\mathbf{u}\cdot\mathbf{n}\,\mathrm{d}\ell,$ (14)
and the corresponding 2D reduction is obtained per unit depth by evaluating the normal flux through the channel walls. Linearity guarantees that the projection structure onto $\phi(x)$ carries over to arbitrary cross-sectional shapes through the corresponding conductance $\kappa(x)$.
We have presented a minimal hydrodynamic theory linking distributed wall exchange to net transport. By projecting wall exchange onto a geometry-dependent sensitivity profile, we show that confinement acts not as a purely resistive barrier but as a geometric rectifier.
The exchange-driven branch exhibits four defining signatures relative to classical Poiseuille flow: (i) viscosity independence at leading order (for prescribed $s$); (ii) inverted confinement scaling, $u\propto r^{-1}$ rather than $r^{2}$; (iii) length amplification, $Q\propto L$ rather than $L^{-1}$ at fixed $\Delta P$; and (iv) linear superposition with pressure-driven throughput.
This rectification framework is not restricted to mass exchange, but applies generically to transport driven by the divergence of boundary currents. If lateral gradients in chemical potential, charge density, or temperature generate a nonuniform slip velocity $u_{s}(x)$, the axial flux takes the form $q(x)=-\kappa(x)\,\mathrm{d}P/\mathrm{d}x+A(x)u_{s}(x)$. Imposing $\mathrm{d}q/\mathrm{d}x=0$ yields an effective source $s(x)\equiv-\frac{\mathrm{d}}{\mathrm{d}x}[A(x)u_{s}(x)]$. Any inhomogeneous surface activity entering through such a boundary-driven flux therefore projects onto the same geometric sensitivity $\phi(x)$ and inherits the defining signatures of exchange-driven transport: leading-order viscosity independence (for prescribed activity), inverted confinement scaling, length amplification, and linear superposition with pressure-driven flow. The conduit thus acts as a universal geometric rectifier, converting spatial variations in surface properties into net axial transport.
In biological settings, walls may be compliant. Since the viscosity independence and the principal scalings arise from the ratio of $\kappa(x)$ to $\mathcal{R}$, the rectification logic remains robust to steady deformation: compliance may alter the detailed form of $A(x)$, but the qualitative “narrower-is-faster” and “longer-is-stronger” trends persist provided geometric asymmetry is maintained.
Finally, our results suggest a multiscale perspective on microvascular hemodynamics. Spatially structured transvascular exchange (Starling filtration) is ubiquitous in microvessels [Starling1896, MichelCurry1999, Woodcock2012, Michel2020], and is not hydrodynamically neutral when coupled to geometric asymmetry [Li2023]. More broadly, microvascular walls exhibit diverse forms of surface activity, including electrokinetic effects associated with the charged glycocalyx/hydration layer [Weinbaum2007, Tarbell2014, Pollack2013FourthPhase] and osmotic/concentration-gradient driven interfacial stresses [Anderson1989], which can enter the axial flux through an effective boundary-driven contribution. Our framework suggests that any such structured activity can be rectified through the same sensitivity profile $\phi(x)$.
In this setting, the geometry–exchange coupling can play two distinct roles. First, it can generate a small but directed baseline throughput via Eq. (9). Second, under strong confinement where transport is controlled by thin lubricating layers, distributed wall exchange can act as an effective lubrication mechanism and reduce the hydraulic load required to sustain perfusion [Li2023]. These predictions motivate quantitative tests in microvessel geometries with controlled exchange patterns.
Acknowledgements.
Acknowledgments to be added.